# Sums of squares of integers from residue classes

### v1

@@ Abstract.
A subset $\mathcal{A}\subseteq\mathbb{Z}$ is called $s$-almost square universal if every sufficiently large positive integer can be written as a sum of at most $s$ squares of integers from $\mathcal{A}$. In this article, we study the minimal number $\mathrm{ASU}(\mathcal{A}_{d,m})$ with this property, where $\mathcal{A}_{d,m}$ denotes the residue class of $d$ modulo $m$, with $m\in\mathbb{N}$ and $d\in\mathbb{Z}$. We further prove that $\mathcal{A}_{d,m}$ is $s$-square universal for some $s\in\mathbb{N}$ if and only if $d\equiv\pm 1\pmod{m}$, and determine the minimal such number $\mathrm{SU}(\mathcal{A}_{d,m})$ in these cases.
@@ 1. Introduction
Waring’s problem asks whether, for a given integer $k\geq 2$, there exists a number $r$ such that every positive integer is a sum of at most $r$ $k$-th powers of positive integers. In 1770, the same year that Edward Waring posed his conjecture, Lagrange proved the Four Square Theorem, which states that every positive integer is the sum of at most four squares of integers. This theorem, though proved independently, resolved Waring’s problem in the special case $k=2$. In general, the existence of such a number was proved by Hilbert [12] in 1909, and the determination of the smallest such number $g(k)$ has been extensively studied.
@@
$\mathrm{SU}({\mathbb{Z}}\setminus 2{\mathbb{Z}})=10,\ \mathrm{SU}({\mathbb{Z}}\setminus 3{\mathbb{Z}})=6,\ \mathrm{SU}({\mathbb{Z}}\setminus 5{\mathbb{Z}})=5,\text{ and }\mathrm{SU}({\mathbb{Z}}\setminus 2{\mathbb{Z}})=4\text{ for any prime }p\geq 7.$
@@ 2.1. Quadratic spaces and lattices
A quadratic $\mathbb{Z}$}-lattice on $V$ is a free $\mathbb{Z}$}-module $L={\mathbb{Z}}e_{1}+\cdots+{\mathbb{Z}}e_{r}$ of rank $r$ such that ${\mathbb{Q}}L=V$. The $r\times r$ symmetric matrix $M_{L}:=B(e_{i},e_{j})$ is called the Gram matrix of $L$ with respect to a basis $e_{1},\ldots,e_{r}$, and we write $L\cong M_{L}$. If the Gram matrix $M_{L}$ is diagonal, we simply write $L\cong\langle Q(e_{1}),\ldots,Q(e_{r})\rangle$. The quadratic form associated to $L$ is defined by $Q_{L}(x_{1},\ldots,x_{r}):=\bm{x}^{t}M_{L}\bm{x}$ where $\bm{x}=(x_{1},\ldots,x_{r})^{t}$. The discriminant $d_{L}$ of $L$ is the determinant of the Gram matrix $M_{L}$, and the level $N_{L}$ is defined to be the smallest positive integer $N$ such that $NM_{L}^{-1}$ has coefficients in $\mathbb{Z}$}. We define the localization of $V$ and $L$ at any prime spot $p$ by $V_{p}:=V\otimes_{\mathbb{Q}}{\mathbb{Q}}_{p}$ and $L_{p}=L\otimes_{\mathbb{Z}}{\mathbb{Z}}_{p}$.
@@
$a\equiv b\ (\mathrm{mod}\ 2),\quad 4a-b^{2}\geq 0\quad\text{and}\quad 4a-b^{2}\neq 4^{\alpha}(8\beta+7)\text{ for some $\alpha,\beta\in{\mathbb{Z}}$}.$ (2.2)
@@
Since the case when $4a-b^{2}>0$ has been proved in [14, Lemma 4.2], let us assume that $4a-b^{2}=0$. First, $b$ should be even since $4a=b^{2}$. Hence $a$ is also even as $a\equiv b\ (\mathrm{mod}\ 2)$. Writing $a=2a^{\prime}$ and $b=2b^{\prime}$, we have $2a^{\prime}=(b^{\prime})^{2}$, so $b^{\prime}=2b^{\prime\prime}$ for some $b^{\prime\prime}\in{\mathbb{Z}}$ and $a=(b^{\prime\prime})^{2}$. Thus, $x_{i}=(b^{\prime\prime})^{2}$ is a solution to (2.1). ∎
@@ 3. Almost square universality of ${\mathcal{A}}_{d,m}$
    Every $n\geq N$ can be expressed as a sum of at most $Mm+3$ squares for ${\mathcal{A}}_{d,m}$, where
    $N=\begin{cases}\frac{25}{4}m^{4}+Md^{2}&\text{if }m\equiv 1\ (\mathrm{mod}\ 2),\\ \frac{m^{4}}{16}+Md^{2}&\text{if }m\equiv 2\ (\mathrm{mod}\ 4),\\ \frac{m^{4}}{4}+Md^{2}&\text{if }m\equiv 0\ (\mathrm{mod}\ 4).\end{cases}$
@@
    $s\geq\begin{cases}m+3&\text{if }m\equiv 1\ (\mathrm{mod}\ 2),\\ Mm+2&\text{if }m\equiv 0\ (\mathrm{mod}\ 2).\end{cases}$
@@
(2) For $n\geq N$, we will show that there exist integers $a,b\in{\mathbb{Z}}$ and $4\leq r\leq Mm+3$ satisfying
@@
Then, since $a$ and $b$ satisfy (2.2), by Lemma 2.1 there exist $x_{1},x_{2},x_{3},x_{4}\in{\mathbb{Z}}$ such that $\sum_{i=1}^{4}x_{i}^{2}=a$ and $\sum_{i=1}^{4}x_{i}=b$. Hence setting $x_{5}=\cdots=x_{r}=0$, equations (3.1) and (3.2) show that $n=\sum_{i=1}^{r}(mx_{i}+d)^{2}$.
We choose $r\in{\mathbb{N}}$ to be the unique $4\leq r\leq Mm+3$ satisfying $t\equiv\frac{r-r_{0}}{m}\ (\mathrm{mod}\ M)$, which is necessary by part (1) of the proposition. To explain the choice of $a$ and $b$, for simplicity, let us set
@@
(Case 1) $m\equiv 1\ (\mathrm{mod}\ 2)$. Note that if $(a,b)=(a_{0},b_{0})$ is an arbitrary integer solution to $\tilde{t}=ma+2db$, then all integer solutions $(a,b)$ to this equation are given by
$(a_{k},b_{k})=(a_{0}-2dk,b_{0}+mk)\text{ with }k\in{\mathbb{Z}}.$
We claim that $(a,b)=(a_{k},b_{k})$ satisfies the second and third conditions in (3.3) for some $0\leq k\leq 9$. If this claim holds, then we may find $(a,b)$ satisfying (3.3) as follows: since $a=\frac{\tilde{t}-2db}{m}$, we have
@@
Since $n=mt+r_{0}d^{2}=m\tilde{t}+rd^{2}\geq\frac{25}{4}m^{4}+Md^{2}$, we have
$\tilde{t}\geq\frac{25}{4}m^{4}+\frac{Mm-r}{m}d^{2}\geq\frac{25}{4}m^{4}-\frac{4d^{2}}{m},$
which implies that the length of the interval in (3.4) is at least $10m$. Therefore, there are at least ten consecutive solutions $(a,b)$ to $\tilde{t}=ma+2db$ with $4a-b^{2}\geq 0$. Among these, one can find a pair $(a,b)=(a_{k},b_{k})$ that satisfies (3.3), as claimed.
Now we prove the claim. Let $i\in\{0,1\}$ be such that $a_{i}\equiv b_{i}\ (\mathrm{mod}\ 2)$, and note that for all $j\in{\mathbb{Z}}$, we have $a_{i+2j}\equiv b_{i+2j}\ (\mathrm{mod}\ 2)$. Define $D_{k}=4a_{k}-b_{k}^{2}$. If $a_{i}\equiv b_{i}\equiv 1\ (\mathrm{mod}\ 2)$, then $D_{i}\equiv 3\ (\mathrm{mod}\ 8)$, so we may take $k=i$. Assume instead that $a_{i}\equiv b_{i}\equiv 0\ (\mathrm{mod}\ 2)$ and write $a_{i}=2a_{i}^{\prime}$ and $b_{i}=2b_{i}^{\prime}$. Then $D_{i}\equiv 0\ (\mathrm{mod}\ 4)$, and we have
@@
In these cases, we may also take $k=i$. The remaining cases are when $b_{i}^{\prime}\equiv 1\ (\mathrm{mod}\ 2)$ with $a_{i}^{\prime}\equiv 0\ (\mathrm{mod}\ 4)$, and when $b_{i}^{\prime}\equiv 0\ (\mathrm{mod}\ 2)$ with $a_{i}^{\prime}\equiv 0\ (\mathrm{mod}\ 2)$. In the former case, note that $D_{i}/4\equiv 7\ (\mathrm{mod}\ 8)$. Since
$\frac{D_{i+2j}}{4}=2a_{i+2j}^{\prime}-b_{i+2j}^{\prime 2}=\frac{D_{i}}{4}-\left[(4d+2b_{i}^{\prime}m)j+m^{2}j^{2}\right]$
and $4d+2b_{i}^{\prime}m\equiv 2\ (\mathrm{mod}\ 4)$, we have $D_{i+2j_{0}}/4\equiv 3\ (\mathrm{mod}\ 8)$ for $j_{0}\in\{1,3\}$ with $4d+2b_{i}^{\prime}m\equiv 2j_{0}\ (\mathrm{mod}\ 8)$. Thus, we may take $k=i+2j_{0}$. For the latter case, since $b_{i+2}^{\prime}\equiv 1\ (\mathrm{mod}\ 2)$, repeating the above argument by replacing $i$ with $i+2$ shows that $(a_{k},b_{k})$ for some $k\in\{i+2,i+4,i+8\}$ satisfies (3.3). This completes the proof of the claim.
(Case 2) $m\equiv 0\ (\mathrm{mod}\ 2)$. The strategy is the same as in the previous case with appropriate modifications. Note that $\tilde{t}\equiv 0\ (\mathrm{mod}\ M)$ and $d\equiv 1\ (\mathrm{mod}\ 2)$. Suppose $(a,b)=(a_{0},b_{0})$ is an arbitrary integer solution to $\frac{\tilde{t}}{2}=\frac{m}{2}a+db$. Then all integer solutions $(a,b)$ are given by
$(a_{k},b_{k})=(a_{0}-dk,b_{0}+\frac{m}{2}k)\text{ with }k\in{\mathbb{Z}}.$
We claim that there exists some integer $0\leq k\leq M-1$ such that $(a,b)=(a_{k},b_{k})$ satisfies the second and third conditions in (3.3).
If $m\equiv 2\ (\mathrm{mod}\ 4)$, then $\frac{\tilde{t}}{2}\equiv 0\ (\mathrm{mod}\ 2)$. Hence for all $j$, we have $a_{j}\equiv b_{j}\ (\mathrm{mod}\ 2)$, and there exists some $i\in\{0,1\}$ such that $a_{i}\equiv b_{i}\equiv 1\ (\mathrm{mod}\ 2)$. For this $i$, it follows that $4a_{i}-b_{i}^{2}\equiv 3\ (\mathrm{mod}\ 8)$, so we may take $k=i$.
If $m\equiv 0\ (\mathrm{mod}\ 4)$, then let $i\in\{0,1\}$ be such that $a_{i+2j}\equiv b_{i+2j}\ (\mathrm{mod}\ 2)$ for all integers $j$. If either $a_{i}\equiv b_{i}\equiv 1\ (\mathrm{mod}\ 2)$, or $a_{i}=2a_{i}^{\prime}$ and $b_{i}=2b_{i}^{\prime}$ satisfy one of the cases in (3.5), then we may take $k=i$. Otherwise, we consider two remaining cases: either $b_{i}^{\prime}\equiv 1\ (\mathrm{mod}\ 2)$ with $a_{i}^{\prime}\equiv 0\ (\mathrm{mod}\ 4)$; or $b_{i}^{\prime}\equiv 0\ (\mathrm{mod}\ 2)$ with $a_{i}^{\prime}\equiv 0\ (\mathrm{mod}\ 2)$. In the first case, since $D_{i}/4\equiv 7\ (\mathrm{mod}\ 8)$, we have
@@
In the second case, we have $a_{i+2}^{\prime}=a_{i}^{\prime}-d\equiv 1\ (\mathrm{mod}\ 2)$ and $b_{i+2}^{\prime}=b_{i}^{\prime}+\frac{m}{2}\equiv 0\ (\mathrm{mod}\ 2)$. Therefore, we have $D_{i+2}/4=2a_{i+2}^{\prime}-b_{i+2^{\prime}}^{2}\equiv 2\ (\mathrm{mod}\ 4)$. Thus, in either case, we may take $k=i+2$. This completes the proof of the claim.
Note that from $n=m\tilde{t}+rd^{2}\geq\frac{m^{4}}{M^{2}}+Md^{2}$, we obtain
$\tilde{t}\geq\frac{m^{4}}{M^{2}}+\frac{Mm-r}{m}d^{2}>\frac{m^{4}}{M^{2}}-\frac{4d^{2}}{m}.$
Hence the length of the interval in (3.4) is greater than $M\cdot\frac{m}{2}$. Therefore, there are at least $M$ consecutive solutions $(a,b)$ to $\frac{\tilde{t}}{2}=\frac{m}{2}a+db$ with $4a-b^{2}>0$, among which one can find $(a,b)=(a_{k},b_{k})$ satisfying (3.3) by the claim. This completes the proof of part (2).
(3) First, suppose $m\equiv 0\ (\mathrm{mod}\ 2)$. Let $p$ be a prime such that $p\equiv 3\ (\mathrm{mod}\ 4)$ and $(p,mM)=1$. Consider a positive integer $n=mMt_{0}+2d^{2}$, where $t_{0}$ is a positive integer such that

### v2

@@ Abstract.
A subset $\mathcal{A}\subseteq\mathbb{Z}$ is called $s$-almost square universal if every sufficiently large positive integer can be written as a sum of at most $s$ squares of integers from $\mathcal{A}$. In this article, we study the minimal number $\mathrm{ASU}(\mathcal{A}_{d,m})$ with this property, where ${\mathcal{A}}_{d,m}$ denotes the residue class of $d$ modulo $m$, with $m\in{\mathbb{N}}$ and $d\in\mathbb{Z}$. We further prove that $\mathcal{A}_{d,m}$ is $s$-square universal for some $s\in\mathbb{N}$ if and only if $d\equiv\pm 1\ (\mathrm{mod}\ m)$, and determine the minimal such number $\mathrm{SU}(\mathcal{A}_{d,m})$ in these cases.
@@ 1. Introduction
Waring’s problem asks whether, for a given integer $k\geq 2$, there exists a number $r$ such that every positive integer is a sum of at most $r$ $k$-th powers of positive integers. In 1770, the same year that Edward Waring posed his conjecture, Lagrange proved the Four Square Theorem, which states that every positive integer is a sum of at most four squares of integers. This theorem, though proved independently, resolved Waring’s problem in the special case $k=2$. In general, the existence of such a number was proved by Hilbert [12] in 1909, and the determination of the smallest such number $g(k)$ has been extensively studied.
@@
$\mathrm{SU}({\mathbb{Z}}\setminus 2{\mathbb{Z}})=10,\ \mathrm{SU}({\mathbb{Z}}\setminus 3{\mathbb{Z}})=6,\ \mathrm{SU}({\mathbb{Z}}\setminus 5{\mathbb{Z}})=5,\text{ and }\mathrm{SU}({\mathbb{Z}}\setminus p{\mathbb{Z}})=4\text{ for any prime }p\geq 7.$
@@ 2.1. Quadratic spaces and lattices
A quadratic $\mathbb{Z}$}-lattice on $V$ is a free $\mathbb{Z}$}-module $L={\mathbb{Z}}e_{1}+\cdots+{\mathbb{Z}}e_{r}$ of rank $r$ such that ${\mathbb{Q}}L=V$. The $r\times r$ symmetric matrix $M_{L}:=B(e_{i},e_{j})$ is called the Gram matrix of $L$ with respect to the basis $e_{1},\ldots,e_{r}$, and we write $L\cong M_{L}$. If the Gram matrix $M_{L}$ is diagonal, we simply write $L\cong\langle Q(e_{1}),\ldots,Q(e_{r})\rangle$. The quadratic form associated to $L$ is defined by $Q_{L}(x_{1},\ldots,x_{r}):=\bm{x}^{t}M_{L}\bm{x}$ where $\bm{x}=(x_{1},\ldots,x_{r})^{t}$. The discriminant $d_{L}$ of $L$ is the determinant of the Gram matrix $M_{L}$, and the level $N_{L}$ is defined to be the smallest positive integer $N$ such that $NM_{L}^{-1}$ has coefficients in $\mathbb{Z}$}. We define the localization of $V$ and $L$ at any prime spot $p$ by $V_{p}:=V\otimes_{\mathbb{Q}}{\mathbb{Q}}_{p}$ and $L_{p}=L\otimes_{\mathbb{Z}}{\mathbb{Z}}_{p}$.
@@
$a\equiv b\ (\mathrm{mod}\ 2),\quad 4a-b^{2}\geq 0\quad\text{and}\quad 4a-b^{2}\neq 4^{\alpha}(8\beta+7)\text{ for any $\alpha,\beta\in{\mathbb{Z}}$}.$ (2.2)
@@
Since the case when $4a-b^{2}>0$ has been proved in [14, Lemma 4.2], let us assume that $4a-b^{2}=0$. First, $b$ should be even since $4a=b^{2}$. Hence $a$ is also even as $a\equiv b\ (\mathrm{mod}\ 2)$. Writing $a=2a^{\prime}$ and $b=2b^{\prime}$, we have $2a^{\prime}=(b^{\prime})^{2}$, so $b^{\prime}=2b^{\prime\prime}$ for some $b^{\prime\prime}\in{\mathbb{Z}}$ and $a=4(b^{\prime\prime})^{2}$. Thus, $x_{i}=b^{\prime\prime}$ is a solution to (2.1). ∎
Lemma 2.2.
For $a,b\in{\mathbb{Z}}$, the system of equations
$\begin{cases}x_{1}^{2}+x_{2}^{2}+x_{3}^{2}+x_{4}^{2}+x_{5}^{2}=a\\ x_{1}+x_{2}+x_{3}+x_{4}+x_{5}=b\end{cases}$ (2.3)
has a solution $x_{1},x_{2},x_{3},x_{4},x_{5}\in{\mathbb{Z}}$ if and only if $a\equiv b\ (\mathrm{mod}\ 2)$ and $5a-b^{2}\geq 0$.
Proof.
The “only if” part follows from the fact that $x_{i}^{2}\equiv x_{i}\ (\mathrm{mod}\ 2)$ and the Cauchy–Schwartz inequality. Hence it suffices to prove the “if” part. If $5a=b^{2}$, then $b=5b^{\prime}$ for some $b^{\prime}\in{\mathbb{Z}}$, and hence $a=5b^{\prime 2}$. In this case, setting $x_{i}=b^{\prime}$ for all $i$ gives a solution. We therefore assume that $5a-b^{2}>0$.
Let $\ell={\mathbb{Z}}v_{1}+{\mathbb{Z}}v_{2}\cong\left(\begin{smallmatrix}5&b\\ b&a\end{smallmatrix}\right)$ and $L={\mathbb{Z}}e_{1}+\cdots+{\mathbb{Z}}e_{5}\cong\langle 1,1,1,1,1\rangle$ be $\mathbb{Z}$}-lattices. Then $\ell$ is locally represented by $L$ since their ranks differ by $3$ and $L_{p}$ is unimodular for every prime $p$. Thus $\ell$ is globally represented by $L$ as the class number of $L$ is one. Hence there exists a representation $\sigma:\ell{\,\rightarrow\,}L$.
By changing the sign of the basis vectors $e_{i}$ and reordering them if necessary, we may assume that
$\sigma(v_{1})=e_{1}+e_{2}+e_{3}+e_{4}+e_{5}\quad\text{ or }\quad\sigma(v_{1})=2e_{1}+e_{5}.$
Write $\sigma(v_{2})=\sum_{i=1}^{5}y_{i}e_{i}$ for some $y_{i}\in{\mathbb{Z}}$. Then $\sum_{i=1}^{5}y_{i}^{2}=B(\sigma(v_{2}),\sigma(v_{2}))=B(v_{2},v_{2})=a$. In the former case, we have $\sum_{i=1}^{5}y_{i}=B(\sigma(v_{1}),\sigma(v_{2}))=B(v_{1},v_{2})=b$, and hence $x_{i}=y_{i}$ gives a solution.
In the latter case, consider the sublattice $K={\mathbb{Z}}w_{1}+{\mathbb{Z}}w_{2}\subseteq L$, where $w_{1}=2e_{1}$ and $w_{2}=y_{1}e_{1}+y_{2}e_{2}+y_{3}e_{3}+y_{4}e_{4}$. Since $K\cong\left(\begin{smallmatrix}4&b-y_{5}\\ b-y_{5}&a-y_{5}^{2}\end{smallmatrix}\right)$ is a sublattice of $L$, it is positive semi-definite. Hence its discriminant $d_{K}=4(a-y_{5}^{2})-(b-y_{5})^{2}\geq 0$. Since $K\subseteq{\mathbb{Z}}e_{1}+\cdots+{\mathbb{Z}}e_{4}\cong\langle 1,1,1,1\rangle$ in addition, $d_{K}$ is not of the form $4^{\alpha}(8\beta+7)$ for any $\alpha,\beta\in{\mathbb{Z}}$. Finally, as $a-y_{5}^{2}\equiv b-y_{5}\ (\mathrm{mod}\ 2)$, Lemma 2.1 yields $x_{1},x_{2},x_{3},x_{4}\in{\mathbb{Z}}$ satisfying
$\begin{cases}x_{1}^{2}+x_{2}^{2}+x_{3}^{2}+x_{4}^{2}=a-y_{5}^{2}\\ x_{1}+x_{2}+x_{3}+x_{4}=b-y_{5}.\end{cases}$
Setting $x_{5}=y_{5}$ then gives a solution. This completes the proof of the lemma. ∎
@@ 3. Almost square universality of ${\mathcal{A}}_{d,m}$
    For $m\equiv 1\ (\mathrm{mod}\ 2)$, every $n\geq\frac{m^{4}}{5}+md^{2}$ is a sum of at most $m+4$ squares in ${\mathcal{A}}_{d,m}^{2}$.
  3. (3)
    For $m\equiv 0\ (\mathrm{mod}\ 2)$, every $n\geq\frac{m^{4}}{M^{2}}+Mmd^{2}$ is a sum of at most $Mm+3$ squares for ${\mathcal{A}}_{d,m}^{2}$.
@@
    $s\geq\begin{cases}m+4&\text{if }m\equiv 1\ (\mathrm{mod}\ 2)\text{ with }m\geq 5,\\ Mm+2&\text{if }m\equiv 0\ (\mathrm{mod}\ 2).\end{cases}$
@@
(2) We choose $r\in{\mathbb{N}}$ to be the unique integer satisfying $5\leq r\leq m+4$ and $r\equiv r_{0}\ (\mathrm{mod}\ m)$, as required by part (1) of the proposition. For $n\geq N$, we claim that there exist integers $a,b\in{\mathbb{Z}}$ satisfying
@@
Once we prove the claim, by Lemma 2.2, there exist $x_{1},x_{2},x_{3},x_{4},x_{5}\in{\mathbb{Z}}$ such that $\sum_{i=1}^{5}x_{i}^{2}=a$ and $\sum_{i=1}^{5}x_{i}=b$. Setting $x_{6}=\cdots=x_{r}=0$, equations (3.1) and (3.2) then show that $n=\sum_{i=1}^{r}(mx_{i}+d)^{2}$.
To prove the claim, for simplicity, set
@@
Note that if $(a,b)=(a_{0},b_{0})$ is an arbitrary integer solution to $\tilde{t}=ma+2db$, then all integer solutions $(a,b)$ to this equation are given by
$(a_{k},b_{k})=(a_{0}-2dk,b_{0}+mk),\quad k\in{\mathbb{Z}}.$
Moreover,
$5a-b^{2}=5\cdot\frac{\tilde{t}-2db}{m}-b^{2}\geq 0$ $\iff mb^{2}+10db-5\tilde{t}\leq 0$
$\iff b\in\left[-\frac{5d}{m}-\frac{\sqrt{25d^{2}+5m\tilde{t}}}{m},-\frac{5d}{m}+\frac{\sqrt{25d^{2}+5m\tilde{t}}}{m}\right].$ (3.4)
Since $n=mt+r_{0}d^{2}=m\tilde{t}+rd^{2}\geq\frac{1}{5}m^{4}+md^{2}$, we obtain
$5m\tilde{t}\geq m^{4}+5(m-r)d^{2}\geq m^{4}-25d^{2}.$
This implies that the length of the interval in (3.6) is at least $2m$. Therefore, there exist at least two consecutive integer solutions $(a,b)$ of $\tilde{t}=ma+2db$ satisfying $5a-b^{2}\geq 0$, one of which necessarily satisfies $a\equiv b\ (\mathrm{mod}\ 2)$. Hence, one can find a pair $(a,b)=(a_{k},b_{k})$ that satisfies (3.3), as claimed.
(3) For $n\geq N$, we show that there exist integers $a,b\in{\mathbb{Z}}$ and an integer $4\leq r\leq Mm+3$ such that
[…]
Since $a$ and $b$ satisfy (2.2), Lemma 2.1 implies that there exist $x_{1},x_{2},x_{3},x_{4}\in{\mathbb{Z}}$ such that $\sum_{i=1}^{4}x_{i}^{2}=a$ and $\sum_{i=1}^{4}x_{i}=b$. Setting $x_{5}=\cdots=x_{r}=0$, equations (3.1) and (3.2) then show that $n=\sum_{i=1}^{r}(mx_{i}+d)^{2}$.
We choose $r\in{\mathbb{N}}$ to be the unique integer with $4\leq r\leq Mm+3$ satisfying $t\equiv\frac{r-r_{0}}{m}\ (\mathrm{mod}\ M)$, as required by part (1) of the proposition. To explain the choice of $a$ and $b$, as before, set
$\tilde{t}:=t-\frac{r-r_{0}}{m}d^{2}.$
Note that $\tilde{t}\equiv 0\ (\mathrm{mod}\ M)$ and $d\equiv 1\ (\mathrm{mod}\ 2)$. Suppose that $(a,b)=(a_{0},b_{0})$ is an arbitrary integer solution of $\frac{\tilde{t}}{2}=\frac{m}{2}a+db$. Then all integer solutions $(a,b)$ are given by
$(a_{k},b_{k})=\left(a_{0}-dk,b_{0}+\frac{m}{2}k\right)\text{ with }k\in{\mathbb{Z}}.$
We claim that there exists some integer $0\leq k\leq\frac{8}{M}-1$ such that $(a,b)=(a_{k},b_{k})$ satisfies the second and the last conditions in (3.5).
Assuming the claim, we now construct such a pair $(a,b)$. Since $a=\frac{\tilde{t}-2db}{m}$, we have
@@
Since $n=m\tilde{t}+rd^{2}\geq\frac{m^{4}}{M^{2}}+Mmd^{2}$, we obtain
$m\tilde{t}\geq\frac{m^{4}}{M^{2}}+(Mm-r)d^{2}>\frac{m^{4}}{M^{2}}-4d^{2}.$
Hence, the length of the interval in (3.6) is greater than $\frac{8}{M}\cdot\frac{m}{2}$. Therefore, there are at least $\frac{8}{M}$ consecutive integer solutions $(a,b)$ of $\frac{\tilde{t}}{2}=\frac{m}{2}a+db$ with $4a-b^{2}>0$, among which one can find a pair $(a,b)=(a_{k},b_{k})$ satisfying (3.5) by the claim.
We now prove the claim. If $m\equiv 2\ (\mathrm{mod}\ 4)$, then $\frac{\tilde{t}}{2}\equiv 0\ (\mathrm{mod}\ 2)$. Hence $a_{j}\equiv b_{j}\ (\mathrm{mod}\ 2)$ for all $j$, and there exists some $i\in\{0,1\}$ such that $a_{i}\equiv b_{i}\equiv 1\ (\mathrm{mod}\ 2)$. For this $i$, we have $4a_{i}-b_{i}^{2}\equiv 3\ (\mathrm{mod}\ 8)$, and we may take $k=i$.
If $m\equiv 0\ (\mathrm{mod}\ 4)$, then let $i\in\{0,1\}$ be such that $a_{i+2j}\equiv b_{i+2j}\ (\mathrm{mod}\ 2)$ for all $j\in{\mathbb{Z}}$. Set $D_{k}=4a_{k}-b_{k}^{2}$. If $a_{i}\equiv b_{i}\equiv 1\ (\mathrm{mod}\ 2)$, then $D_{i}\equiv 3\ (\mathrm{mod}\ 8)$, and we may take $k=i$. Otherwise, assume $a_{i}\equiv b_{i}\equiv 0\ (\mathrm{mod}\ 2)$, and write $a_{i}=2a_{i}^{\prime}$ and $b_{i}=2b_{i}^{\prime}$. Then $D_{i}\equiv 0\ (\mathrm{mod}\ 4)$, and
@@
In these cases, we may again take $k=i$. The remaining cases are:
  1. (i)
    $b_{i}^{\prime}\equiv 1\ (\mathrm{mod}\ 2)$ with $a_{i}^{\prime}\equiv 0\ (\mathrm{mod}\ 4)$; and
  2. (ii)
    $b_{i}^{\prime}\equiv 0\ (\mathrm{mod}\ 2)$ with $a_{i}^{\prime}\equiv 0\ (\mathrm{mod}\ 2)$.
In case (i), since $D_{i}/4\equiv 7\ (\mathrm{mod}\ 8)$, we have
@@
In case (ii), we have $a_{i+2}^{\prime}=a_{i}^{\prime}-d\equiv 1\ (\mathrm{mod}\ 2)$ and $b_{i+2}^{\prime}=b_{i}^{\prime}+\frac{m}{2}\equiv 0\ (\mathrm{mod}\ 2)$., and hence $D_{i+2}/4=2a_{i+2}^{\prime}-b_{i+2}^{\prime 2}\equiv 2\ (\mathrm{mod}\ 4)$. Thus, in either case, we may take $k=i+2$. This proves the claim and completes the proof of part (3).
(4) First, suppose that $m\equiv 0\ (\mathrm{mod}\ 2)$. Let $p$ be a prime such that $p\equiv 3\ (\mathrm{mod}\ 4)$ and $(p,mM)=1$. Consider a positive integer $n=mMt_{0}+2d^{2}$, where $t_{0}$ is a positive integer satisfying